# Stability of spherical thin-shell wormholes in scalar-tensor theories

### v1

@@ 4 Wormholes with a radial electric field
In this section, we consider two examples of interest. They correspond to particular cases of Einstein-Maxwell theories coupled to a scalar field. We adopt a solution from each theory to construct the corresponding thin-shell wormholes and we analyze the stability of the static configurations and the fulfillment of the WEC.
4.1 Constant scalar field
@@
In order to obtain a traversable wormhole, the throat radius should satisfy $a_{0}>r_{h}$ when $Q\leq Q_{c}$, which removes the horizon and the singularity at the origin, and $a_{0}>0$ if $Q>Q_{c}$, which takes away the naked singularity at the origin. We use the formalism of the previous section to analyze the behavior of the static configurations. The possible values of $a_{0}$ are those which satisfy Eq. (31) and the stability is determined by Eq. (30). The scalar field shown in Eq. (46) should be real, which restricts the value of charge to $Q^{2}<-s$ when $s<0$.
@@
4.2 Variable scalar field
A more general action of the form (1) for Einstein-Maxwell theory with a conformally coupled self interacting scalar field $\phi$ has been presented in [14], with $f(\phi)$ given again by Eq. (36), the electromagnetic Lagrangian by Eq. (38), and
$U(\phi)=\frac{\Lambda}{16\pi G}+\alpha\phi^{4},$ (51)
where $\Lambda$ is the cosmological constant and $\alpha$ is a constant. We notice that this action involves a term of higher order in the scalar field in addition to the one that appears the previous example. The field equations are
$G_{\mu\nu}+\Lambda g_{\mu\nu}=8\pi G\left(T_{\mu\nu}^{\phi}+T_{\mu\nu}^{EM}\right),$ (52)
$\Box\phi-\frac{1}{6}R\phi-4\alpha\phi^{3}=0,$ (53)
and Eq. (41) for the electromagnetic field. The total energy-momentum tensor $T_{\mu\nu}$ is the sum of
$T_{\mu\nu}^{\phi}=\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{2}g_{\mu\nu}g^{\alpha\beta}\partial_{\alpha}\phi\partial_{\beta}\phi+\frac{1}{6}\left[g_{\mu\nu}\Box-\nabla_{\mu}\nabla_{\nu}+G_{\mu\nu}\right]\phi^{2}-\alpha g_{\mu\nu}\phi^{4}$ (54)
and $T_{\mu\nu}^{EM}$ shown in Eq. (43).
We are interested again in a charged spacetime whose metric is solution of Eq. (52), having the metric [14]
$ds^{2}=-\left[-\frac{\Lambda}{3}r^{2}+\left(1-\frac{m}{r}\right)^{2}\right]dt^{2}+\left[-\frac{\Lambda}{3}r^{2}+\left(1-\frac{m}{r}\right)^{2}\right]^{-1}dr^{2}+r^{2}d\Omega^{2}\,$, (55)
where $m$ is the mass. We have again used units that $G=1$ and $\mu_{0}/(4\pi)=1$. It is peculiar that the charge does not appear explicitly in the metric, in contrast to the standard Reissner-Nordström spacetime. Instead, the charge has a fixed value determined by the other parameters. Particularly, the charge-to-mass ratio is
$\left(\frac{Q}{m}\right)^{2}=1+\frac{2\pi\Lambda}{9\alpha},$ (56)
which imposes the following constraint
$\frac{\Lambda}{\alpha}\geq-\frac{9}{2\pi}.$ (57)
Therefore, the quotient $\Lambda/\alpha$ has a lower bound. The scalar field, which is a solution of Eq. (53), takes the following form
$\phi(r)=\sqrt{\frac{3}{4\pi}}\frac{\sqrt{m^{2}-Q^{2}}}{r-m},$ (58)
or, by using Eq. (56),
$\phi(r)=\sqrt{-\frac{\Lambda}{6\alpha}}\;\frac{m}{r-m},$ (59)
which for being real requires that
$\frac{Q^{2}}{m^{2}}\leq 1,\qquad\frac{\Lambda}{\alpha}\leq 0.$ (60)
The only non-vanishing component of the electromagnetic field is given by (44). Considering that $Q/m$ gets a fixed value for a given $\Lambda/\alpha$, it is useful to introduce a new dimensionless parameter $\epsilon=\Lambda/\alpha$. Thus, the previous expressions can be rewritten in the form
$\left(\frac{Q}{m}\right)^{2}=\left(1+\frac{2\pi\epsilon}{9}\right),$ (61)
$\phi(r)=\sqrt{-\frac{\epsilon}{6}}\frac{m}{r-m},$ (62)
with the constraint
$\frac{9}{2\pi}\leq\epsilon\leq 0.$ (63)
It becomes clear that as long as $\Lambda$ and $\alpha$ have opposite signs, we do not have to worry about their relative behavior when we use the parameter $\epsilon$. Then, for our calculations, we can write the final expressions in terms of $\Lambda$ and $\epsilon$. The information about $\alpha$ is not lost since each value of $\Lambda$ corresponds to a specific value of $\alpha$ through the value of $\epsilon$.
We can now proceed to analyze the solution that we will use for constructing the thin-shell wormholes. This solution presents two possible cases: one with $\Lambda\geq 0$, which describes a static and spherically symmetric black hole, and the other with $\Lambda<0$, which corresponds to a naked singularity. For $\Lambda>0$, there are an inner (Cauchy), an event, and a cosmological horizons, given by
$r_{i}=\frac{\sqrt{3}}{2\sqrt{\Lambda}}\left(-1+\sqrt{1+\frac{4m\sqrt{\Lambda},}{\sqrt{3}}}\right),$ (64)
$r_{h}=\frac{\sqrt{3}}{2\sqrt{\Lambda}}\left(1-\sqrt{1-\frac{4m\sqrt{\Lambda},}{\sqrt{3}}}\right),$ (65)
and
$r_{c}=\frac{\sqrt{3}}{2\sqrt{\Lambda}}\left(1+\sqrt{1-\frac{4m\sqrt{\Lambda}}{\sqrt{3}}}\right),$ (66)
respectively, which satisfy $r_{i}\leq r_{h}\leq r_{c}$. By computing the Kretschmann invariant
$\mathcal{K}=\frac{48m^{2}(r-m)^{2}}{r^{8}},$ (67)
we conclude that the solution presents a singularity at the origin. The horizons $r_{h}$ and $r_{c}$ both exist when $0<\Lambda m^{2}\leq 3/16$, they coincide if $\Lambda m^{2}=3/16$, while for $\Lambda m^{2}>3/16$ they vanish but the inner horizon $r_{i}$ remains. When $\Lambda<0$ the metric function is always positive and thus, there are no horizons and the singularity is naked.
Figure 2: Stability of a spherical thin-shell wormholes symmetric across the throat, in Einstein-Maxwell theory with a conformally coupled and variable scalar field, in the case of cosmological constant $\Lambda>0$ for different values of the parameter $\epsilon=\Lambda/\alpha$, which determines the squared charge $Q^{2}/m^{2}$ divided by the mass $m$. The dashed line represents the unstable solutions with radius $a_{0}/m$ while there are no stable configurations. The dark gray areas have no physical meaning (see text). The shaded areas represent the regions where the WEC is satisfied.
Figure 3: Stability of a spherical thin-shell wormholes symmetric across the throat, in Einstein-Maxwell theory with a conformally coupled and variable scalar field, in case of cosmological constant $\Lambda<0$ for different values of $\epsilon=\Lambda/\alpha$, which determines the squared charge $Q^{2}/m^{2}$ divided by the mass $m$. The solid line represents static stable configurations and the dashed line shows the unstable solutions, with radius $a_{0}/m$. The shaded areas represent the regions where the WEC is satisfied.
In our construction, when $0<\Lambda m^{2}<3/16$ we require that the throat radius should satisfy $r_{h}<a_{0}<r_{c}$, to avoid the presence of the event horizon, and both the Cauchy horizon and the singularity inside it; in this case, we have a traversable wormhole with a cosmological horizon at each side of the throat. If $\Lambda m^{2}>3/16$ the only horizon is $r_{i}$, with the metric function $A(r)$ positive for $r<r_{i}$ and negative for $r>r_{i}$. This suggests that $r_{i}$ would behave as a kind of cosmological horizon, which surrounds a naked singularity at the origin; it is not clear if the attribute of being a Cauchy horizon is maintained or not in this case. We adopt $0<a_{0}<r_{i}$ to remove the singularity and also to have a positive metric function. When $\Lambda<0$ we only require that $a_{0}>0$ so the singularity at the origin is not present. We use the formalism of the previous section, so $a_{0}$ has to fulfill Eq. (31), while the stability analysis comes from Eq. (30). The scalar field shown in Eqs. (58) and (59) should be real, which imposes the restrictions (60). It is important to emphasize that the parameter $\epsilon$ does not affect the junction condition (26) nor the stability, determined by the second derivative of the potential (30). These equations involve only the metric function, which only depends on the mass and the cosmological constant. The WEC is satisfied when both inequalities in Eq. (35) hold. The parameter $\epsilon$ determines the WEC because it involves the scalar field in which the parameter $\alpha$ appears.
We graphically show the results for the case $\Lambda>0$ in Fig. 2. We analyze the existence and the stability of the static configurations, as well as the WEC, by plotting $a_{0}/m$ against $\Lambda m^{2}$ for different values of the parameter $\epsilon$ and consequently of $Q/m$. The black curve corresponds to the solutions that satisfy (31), where the stable ones are drawn with a solid line and the unstable ones with a dashed line. The dark gray areas represent regions without physical meaning while the dashed areas correspond to regions in which the WEC is satisfied. We see that there is one solution, which is unstable for the whole range of the parameters; it has $a_{0}/m>1$ when $0<\Lambda m^{2}<3/16$ and $a_{0}/m<1$ when $\Lambda m^{2}>3/16$. Additionally, the WEC is always violated at the throat if $0<\Lambda m^{2}<3/16$. However, for $\Lambda m^{2}>3/16$ when the absolute value $\epsilon$ is large enough (i. e. small enough $Q^{2}/m^{2}$) the solution satisfies WEC, but at the expense of $f(\phi_{0})<0$.
We continue now with the special case of $\Lambda=0$. The metric takes the form of the extremal Reissner-Nordström spacetime in which $Q=m$; it has a horizon at $r_{h}=m$ and the scalar field vanishes. In this case, there are no real and positive solutions of Eq. (31), then our construction is not possible.
Finally, we present the results for $\Lambda<0$ in Fig. 3, which shows the more representative results. In this case, for $(-107-51\sqrt{17})/512\approx-0.62\leq\Lambda m^{2}<0$ there is a pair of solutions of Eq. (31), the larger one is stable and the other is unstable. Depending on the value of $\epsilon$ and the other parameters these solutions can satisfy the WEC or not. For a small absolute value $|\epsilon|$ (i.e. large $Q^{2}/m^{2}$), the WEC is fulfilled in a short part of the stable curve which grows as $|\epsilon|$ increases to finally satisfy it along the complete line. After that, the unstable solution also begins to satisfy the WEC and for large enough $|\epsilon|$ it always does. Again, the throat can have normal matter but at the cost of having that $f(\phi_{0})<0$.
It is worthy to mention that the WEC is always satisfied in the bulk. The energy-momentum tensor in an orthonormal frame reads
$T^{\hat{\mu}\hat{\nu}}=\frac{m^{2}}{8\pi r^{4}}\mathrm{diag}\left(1,-1,1,1\right),$ (68)
which for a perfect fluid takes the form $T^{\hat{\mu}\hat{\nu}}=\mathrm{diag}(\rho,p_{1},p_{2},p_{3})$, with $\rho$ the volume energy density and $p_{i}$ the pressures; it is easy to check that $\rho\geq 0$ and $\rho+p_{i}\geq 0$, so the WEC is fulfilled.
To sum up, when $\Lambda>0$ we have found one solution which is always unstable; if $\Lambda=0$ our construction is not possible; while for $\Lambda<0$ we have obtained a pair of solutions, one stable and the other unstable. The configurations can satisfy the WEC or not, depending on the particular values of $\epsilon=\Lambda/\alpha$ and $\Lambda m^{2}$, but the presence of normal matter requires $f(\phi_{0})<0$. The consequence, as mentioned above, is a negative effective Newton constant and a graviton with a ghost behavior. The WEC is satisfied at the bulk in all cases.
@@ 5 Conclusions
In this work, we have constructed a class of spherical thin-shell wormholes in scalar-tensor theories with a self-interaction potential non-minimally coupled to gravity, in which the throat is located at the shell. For wormholes symmetric across the throat, we have studied the stability of static configurations under perturbations preserving the spherical symmetry, establishing a condition in terms of the second derivative of a potential.
We have found that the symmetry across the spherical throat implies that $2f^{\prime}(\phi)[K]=n^{\mu}[\partial_{\mu}\phi]=0$ and consequently $[K]=0$, which brought the equations to a similar form as in the $F(R)$ gravity counterpart. Thin-shells wormholes in $F(R)$ theories have been studied extensively in [56], in which the condition $[K]=0$ also holds [64]. The expressions for the energy density $\sigma$, the pressure $p$, and the condition that the throat radius $a$ should satisfy, have a similar form to the corresponding ones in $F(R)$ theories. They differ by a factor proportional to $f(\phi)$ instead of a function of the Ricci scalar $R$. This behavior is not surprising, since $F(R)$ gravity and scalar-tensor theories are closely related. However, the solutions to use in the wormhole construction depend on the theory adopted, so the specific models present different characteristics.
In particular, we have considered two examples of charged thin-shell wormholes with spherical symmetry, by using known black hole solutions in Einstein-Maxwell theory with a conformally self-interacting scalar field $\phi(r)$. We have studied the stability of the static configurations under radial perturbations and analyzed the fulfillment of the WEC at the throat and in the bulk. In the first example, where the Lagrangian of the theory is quadratic in the scalar field, the adopted solution has a constant $\phi$ and a metric depending on the mass $m$, the charge $Q$, and the scalar hair $s$. We have concluded that there is a pair of stable and unstable configurations for a range of values of $s/m^{2}$ and $Q/m$. The stable solutions appear when $0<s/m^{2}<1$ for large $Q>Q_{c}$ and when $1\leq s/m^{2}<9/8$ for comparatively small values of $Q/m$; in both cases with $a_{0}/m$ close to $1$. In the second example, the Lagrangian includes a cosmological constant $\Lambda$ and a term of fourth order in the scalar field, with a coupling parameter $\alpha$. The chosen solution has a variable $\phi(r)$ and a metric depending on $m$ and $\Lambda$; the charge per mass $Q/m$ is determined by $\Lambda/\alpha$. For $\Lambda>0$ we have shown that there exist one unstable solution and when $\Lambda<0$ we have found two solutions, the one with larger $a_{0}/m$ is stable while the other is unstable.
In both examples, we have verified that the existence of exotic matter cannot be avoided at the throat when requiring the physically sound condition that $f(\phi)>0$, since for a wormhole symmetric across the throat the energy conditions are always violated in any scalar-tensor theory, as it has been already proved [63].

### v2

@@ 4 Wormholes with a radial electric field
In this section, we consider a particular case of Einstein-Maxwell theory coupled to a conformally invariant scalar field. We adopt the solution found in [17] to construct the thin-shell wormholes and we analyze the stability of the static configurations and the fulfillment of the WEC.
@@
In order to obtain a traversable wormhole, the throat radius should satisfy $a_{0}>r_{h}$ when $Q\leq Q_{c}$, which removes the horizon and the singularity at the origin, and $a_{0}>0$ if $Q>Q_{c}$, which takes away the naked singularity at the origin. The scalar field shown in Eq. (52) satisfies that $\phi^{\prime}(r)=0$, then we use the formalism of the previous section to analyze the behavior of the static configurations, with the potential given by Eq. (35). The possible values of $a_{0}$ are those which satisfy Eq. (33) and the stability is determined by Eq. (37). The scalar field should be real, which restricts the value of charge to $Q^{2}<-s$ when $s<0$.
@@ 5 Conclusions
In this work, we have constructed a class of spherical thin-shell wormholes in scalar-tensor theories with a self-interaction potential non-minimally coupled to gravity, in which the throat is located at the shell. For wormholes symmetric across the throat, we have studied the stability of the static configurations under perturbations preserving the spherical symmetry, establishing a condition in terms of the second derivative of a potential.
We have found that when $n^{\mu}\partial_{\mu}\phi]=0$, which implies $[K]=0$, the equations are greatly simplified and resemble those in the $F(R)$ gravity counterpart. The condition $[K]=0$ always holds in the junction formalism of $F(R)$ theories [66]. In our case, the expressions for the condition satisfied by the throat radius $a$, the energy density $\sigma$, and the pressure $p$, have a similar form to the corresponding ones for thin-shell wormholes in $F(R)$ theories. They differ by a factor proportional to $f(\phi)$ instead of a function of the Ricci scalar $R$. This behavior is not surprising, since $F(R)$ gravity and scalar-tensor theories are closely related. However, the solutions to use in the wormhole construction depend on the theory adopted, so the specific models present different characteristics.
In particular, we have considered the example of charged thin-shell wormholes with spherical symmetry, by using known black hole solutions [17] in Einstein-Maxwell theory with a conformally self-interacting scalar field $\phi(r)$. We have studied the stability of the static configurations under radial perturbations and analyzed the fulfillment of the WEC at the throat and in the bulk. The Lagrangian of the theory is quadratic in the scalar field, the spacetime adopted has a constant $\phi$ and a metric depending on the mass $m$, the charge $Q$, and the scalar hair $s$. We have concluded that there is a pair of solutions, one stable and the other unstable, for a range of values of $s/m^{2}$ and $Q/m$. The stable configuration appears when $0<s/m^{2}<1$ for large $Q>Q_{c}$ and when $1\leq s/m^{2}<9/8$ for comparatively small values of $Q/m$; in both cases with $a_{0}/m$ close to $1$. We have verified that the existence of exotic matter at the throat cannot be avoided when requiring the physically sound condition that $f(\phi)>0$, since for a wormhole symmetric across the throat the energy conditions are always violated in any scalar-tensor theory, as it has been already proved [65].
For future research, it would be interesting to consider in detail more complex scenarios in which $n^{\mu}\partial_{\mu}\phi]\neq 0$, especially to study the stability and to analyze whether the presence of exotic matter can be eliminated or reduced.